### v1

@@ 4.2 Alternating Backpropagation
Similar to the backpropagation of the ANN, Alternating Backpropagation also use the chain rule to calculate the gradient of each layer of the X-Net. It’s just that Alternating Backpropagation alternately update the activation function and parameters of the neural network during backpropagation. In the process of Alternating Backpropagation, we treat the output of each neurons as a optimizabale variable $\mathcal{E}$, $\mathcal{E}=f(\mathcal{E}_{left},\mathcal{E}_{right})$, where $\mathcal{E}$ is the output of the current node, $\mathcal{E}_{left}$ is the output of the left node of the current node, $\mathcal{E}_{right}$ is the output of the right node of the current node, $f$ is The numerical computation corresponding to the node activate function. First, take the partial derivative of $\mathcal{E}$ with respect to the loss function, and then calculate the partial derivative of the $\mathcal{E}_{left}$ and $\mathcal{E}_{right}$ with respect to the loss function by the chain rule[15]. Finally, update them with the stochastic gradient descent(SGD)[16, 17] (It can also be other numerical optimization algorithms, such as BFGS and so on). Similar to a conventional ANN, we utilize SGD to update the network’s constants,$w_{new}=w-\alpha*\frac{\partial Loss}{\partial w}$, Where $w_{new}$ is the updated constant, $w$ is the constant without updating, $\alpha$ is the learning rate,$\frac{\partial Loss}{\partial w}$ is the partial derivative of the loss function to $w$. For example, for the neural network architecture shown in Fig. 3c, Through the Alternating Backpropagation, we obtain the gradient of each node’s output $\mathcal{E}_{n}$ and its corresponding two constants $w_{n},b_{n}$. Then, we optimize activate function and parameters in alternation. The specific formula is deduced as follows:
$\displaystyle d\mathcal{E}_{1}$ $=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{1}}\vspace{3ex}$ (11)
$dw_{1}$ $=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{1}}\frac{\partial E_{1}}{\partial w_{1}}=d\mathcal{E}_{1}\frac{\partial\mathcal{E}_{1}}{\partial w_{1}}\vspace{3ex}$ (12)
$db_{1}$ $=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{1}}\frac{\partial\mathcal{E}_{1}}{\partial b_{1}}=d\mathcal{E}_{1}\frac{\partial\mathcal{E}_{1}}{\partial b_{1}}=d\mathcal{E}_{1}\vspace{3ex}$ (13)
$d\mathcal{E}_{2}$ $=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{2}}=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{1}}\frac{\partial\mathcal{E}_{1}}{\partial\mathcal{E}_{2}}=d\mathcal{E}_{1}\frac{\partial\mathcal{E}_{1}}{\partial\mathcal{E}_{2}}\vspace{3ex}$ (14)
$dw_{2}$ $=\frac{\partial\mathcal{L}}{\partial w_{2}}=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{1}}\frac{\partial\mathcal{E}_{1}}{\partial\mathcal{E}_{2}}\frac{\partial\mathcal{E}_{2}}{\partial w_{2}}=d\mathcal{E}_{2}\frac{\partial\mathcal{E}_{2}}{\partial w_{2}}\vspace{3ex}$ (15)
$db_{2}$ $=\frac{\partial\mathcal{L}}{\partial b_{2}}=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{1}}\frac{\partial\mathcal{E}_{1}}{\partial\mathcal{E}_{2}}\frac{\partial\mathcal{E}_{2}}{\partial b_{2}}=d\mathcal{E}_{2}\frac{\partial\mathcal{E}_{2}}{\partial b_{2}}=d\mathcal{E}_{2}\vspace{3ex}$ (16)
$d\mathcal{E}_{3}$ $=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{3}}=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{1}}\frac{\partial\mathcal{E}_{1}}{\partial\mathcal{E}_{2}}\frac{\partial\mathcal{E}_{2}}{\partial\mathcal{E}_{3}}=d\mathcal{E}_{2}\frac{\partial\mathcal{E}_{2}}{\partial\mathcal{E}_{3}}\vspace{3ex}$ (17)
$dw_{3}$ $=\frac{\partial\mathcal{L}}{\partial w_{3}}=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{1}}\frac{\partial\mathcal{E}_{1}}{\partial\mathcal{E}_{2}}\frac{\partial\mathcal{E}_{2}}{\partial\mathcal{E}_{3}}\frac{\partial\mathcal{E}_{3}}{\partial w_{3}}=d\mathcal{E}_{3}\frac{\partial\mathcal{E}_{3}}{\partial w_{3}}\vspace{3ex}$ (18)
$db_{3}$ $=\frac{\partial\mathcal{L}}{\partial b_{3}}=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{1}}\frac{\partial\mathcal{E}_{1}}{\partial\mathcal{E}_{2}}\frac{\partial\mathcal{E}_{2}}{\partial\mathcal{E}_{3}}\frac{\partial\mathcal{E}_{3}}{\partial b_{3}}=d\mathcal{E}_{3}\frac{\partial\mathcal{E}_{3}}{\partial b_{3}}=d\mathcal{E}_{3}\vspace{3ex}$ (19)
$\cdots$
Update the parameters:
$w_{1-new}$ $=w_{1}-\alpha dw_{1}\vspace{3ex}$ (20)
$w_{2-new}$ $=w_{2}-\alpha dw_{2}\vspace{3ex}$ (21)
$w_{3-new}$ $=w_{3}-\alpha dw_{3}\vspace{3ex}$ (22)
$b_{1-new}$ $=b_{1}-\alpha db_{1}\vspace{3ex}$ (23)
$b_{2-new}$ $=b_{2}-\alpha db_{2}\vspace{3ex}$ (24)
$b_{3-new}$ $=b_{3}-\alpha db_{3}\vspace{3ex}$ (25)
$\cdots$
$\mathcal{E}_{1-new}$ $=\mathcal{E}_{1}-\alpha d\mathcal{E}_{1}\vspace{3ex}$ (26)
$\mathcal{E}_{2-new}$ $=\mathcal{E}_{2}-\alpha d\mathcal{E}_{2}\vspace{3ex}$ (27)
$\mathcal{E}_{3-new}$ $=\mathcal{E}_{3}-\alpha d\mathcal{E}_{3}\vspace{3ex}$ (28)
$\cdots$
4.3 Selection of Neuronal Types (differential gene expression.)
After backpropagation, we update the output $\mathcal{E}$ of each neuron in the direction of gradient descent. $\mathcal{E}_{i-new}=\mathcal{E}_{i}-\alpha\frac{\partial L}{\partial\mathcal{E}_{i}}$, The updated neuron output is arranged according to the preorder traversal as $\mathcal{E}_{new}=[\mathcal{E}_{1-new},\mathcal{E}_{2-new},..,\mathcal{E}_{n-new}]$, For the updated output value of each neuron, we update the activate function in turn according to the following method. For the $i^{th}$ node, Suppose its symbol is $+$, that is, $\mathcal{E}_{i-old}=\mathcal{E}_{l-old}+\mathcal{E}_{r-old}$, where $\mathcal{E}_{l-old}$ is the output of the left node of the $i^{th}$ node, $\mathcal{E}_{r-old}$ is the output of the right node of the $i^{th}$ node. Assuming that the output value of the $i^{th}$ node is updated to $\mathcal{E}_{i-new}$, and its left and right child nodes are updated to be $\mathcal{E}_{l-new}$ and $\mathcal{E}_{r-new}$, respectively. We denote them as $\mathcal{E}_{left},\mathcal{E}_{right}$, Then these two values are taken as inputs into the activation function library. (the unary activate function uses only the $\mathcal{E}_{left}$ as input), and get the output list: $[\mathcal{E}_{left}+\mathcal{E}_{right},\mathcal{E}_{left}-\mathcal{E}_{right},\mathcal{E}_{left}*\mathcal{E}_{right},\mathcal{E}_{left}\div\mathcal{E}_{right},sin(\mathcal{E}_{left}),cos(\mathcal{E}_{left}),log(\mathcal{E}_{left}),\sqrt{\mathcal{E}_{left}},e^{\mathcal{E}_{left}},relu(x),\\ sigmoid(x),x]$. These output values correspond to the 12 activation functions of [+,,×,÷$[+,-,\times,\div,sin,cos,log,sqrt,exp,relu,sigmoid,x]$. Then we make the difference between these 12 outputs and $\mathcal{E}_{i-new}$ respectively and then take the absolute value to get $G=[g_{1},g_{2},...,g_{1}2]$. Next, we choose the smallest of the $G$, and then, if this value is also less than the preset threshold(e.g.0.01), we use its corresponding activate function as the new activate function for the $i^{th}$ node. Other nodes are analogous.
4.4 Substitution of Activation Functions (Neuronal Evolution)
In order to ensure that the network can still forward propagation normally after the structure and neuron activation function are updated. For instance, the $x$ node can no longer contain child nodes, the binary activate function requires two child nodes, etc. For this purpose we have established the following rules:(1) When the unary activate function becomes a unary activate function, the node symbol is directly transformed, and the others do not make any changes. (2) When the unary activate function is transformed into a binary activate function, the original child node of the unary activate function is retained, and the other child node of None is replaced by x. (3) When the binary activate function becomes a binary activate function, the node symbol is directly changed, and no other changes are made. (4), when the binary activate function is transformed into the unary activate function, there are two situations at this time. a, when it is better to use the output of the left child node as the input, only keep the left child node. When the output of the right child node is used as the input and the result is better, only the right child node is kept. (3), regardless of any activate function, if it becomes x, all nodes whose main node is this node will become None. (4), there are two cases when x becomes an activate function. a, if x becomes a unary activate function, the node symbol becomes a unary activate function, and the left node is added to the leaf node x. When x becomes a binary operator, the node becomes a binary activate function, and the left and right nodes are added, both of which are leaf nodes x. The specific operator replacement process is shown more vividly in Fig. 4.
4.5 Restrictions and handling of numerical computing
To avoid infinite values ”inf” and NaN values during forward and backward propagation, we place some restrictions on computation, (1) To prevent numerical overflow, the output values of all nodes are truncated during the computation process[18], Assuming that the output value of a node is $\mathcal{E}$, we allow the maximum output range to be $[-V,+V]$, then, when $|\mathcal{E}|>=V$, =||*$\mathcal{E}=\frac{\mathcal{E}}{|\mathcal{E}|}*V$ (2) In order to prevent the gradient from disappearing or exploding during the backpropagation of the network, we truncate the gradient and limit the gradient range[19, 20]. Assuming that the value of the gradient is $g$, the output range is$[-G_{max},-G_{min}]\cup[G_{min},G_{max}]$, When $|g|>=G_{max}$ or $|g|<=G_{min}$, The values of $g$ are respectively: $g=\frac{g}{|g|}*G_{max}$ and $g=\frac{g}{|g|}*G_{min}$. (3) We restrict the generation of some unreasonable operations. such as for $log(x)$, where the value of $x$ must be greater than 0, Therefore, operators with negative values in the range cannot be in the position of $x$. For example, $log(sin(x))$, $\sqrt{cos(x)}$ and similar operations are not allowed.
@@ 4.6 Avoid getting stuck in local optimum
In the process of optimization and training, the network is likely to fall into a local optimum if no additional conditions are added[21]. In order to prevent this from happening as much as possible, we set a count value of $count$. Whenever the network loss does not change or increase, the $count$ increases by 1. When the $count$ reaches a certain threshold, we arbitrarily select a node of the neural network, and arbitrarily change its activate function to other functions in the activate function library, By adding such a ”disturbance”, the algorithm can jump out of the local optimum as much as possible. We found through experiments that better results can be obtained when the threshold value is 20.
4.7 Step size adaptation function: Ada-$\alpha$
In the process of network training, the step size has an extremely important influence on the convergence speed of the network and whether the optimal solution can be found[22, 23, 24]. Therefore, in order to better adapt to the algorithm and make the algorithm achieve better results, we designed an Ada-$\alpha$ that can adaptively and dynamically adjust the step size according to the network training situation. The formula for Ada-$\alpha$ is as follows(29):
$\alpha=\frac{tanh(e^{-|Loss[-2]-Loss[-1]|})}{a}.$ (29)
$Loss[-2]$ represents the loss value of the previous iteration, $Loss[-1]$ represents the loss value of the current iteration, and $a$ is the hyperparameter for adjusting the step size range. The Ada-$\alpha$ is inspired by people going down the mountain. For example, when we are descending a relatively steep hill, our bodies will lean back, one small step at a time. And when the mountain is relatively flat, we will boldly take big strides forward. As we were walking, the altitude we were at suddenly increased. This may be because our stride is too big, ”fly” to the other side of the mountain. At this moment, we should take smaller steps to prevent us from striding in the wrong direction. Inspired by this, we propose the Ada-$\alpha$. When the value of the front and rear loss functions decreases greatly, we make the step size $\alpha$ smaller. When the value of the front and rear loss functions decreases relatively small, we make the step size $\alpha$ larger. When the value of the front and rear loss functions rises. That means step size $\alpha$ is too big, and we’re flying to the other side of the mountain, so we’re going to make $\alpha$ smaller.
5 Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant 92370117, in part by CAS Project for Young Scientists in Basic Research under Grant YSBR-090, and in part by the Key Research Program of the Chinese Academy of Sciences under Grant XDPB22
Appendix A Section title of first appendix

### v2

@@ 1.2 Alternating Backpropagation
While traditional backpropagation optimizes a network’s parameters, X-Net requires updating the parameters and activation functions. Therefore, we designed the alternating backpropagation algorithm to update X-Net’s parameters and activation functions alternately. In the process of alternating backpropagation, each neuron is viewed as a variable $E_{i}=f_{i}(E_{\rm left}^{i},E_{\rm right}^{i})$ requiring optimization, where $E_{i}$ is the output of the $i^{\rm th}$ node, $E_{\rm left}^{i}$ and $E_{\rm right}^{i}$ are the left and right inputs of the node $i$, respectively. $f_{i}$ is the numerical computation corresponding to the node activation function. We calculate the derivative of $E_{\rm left}^{i}$ and $E_{\rm right}^{i}$ by the chain rule  (?) and update them with the stochastic gradient descent (SGD)  (?, ?), or other numerical optimization algorithms such as BFGS, L-BFGS. After the gradients have been updated, the activation functions are selected on the basis of $E^{i}_{\rm new}$ (details of selection and substitution are described in section 1.3 and 1.4, respectively), then the constants in the network are updated by SGD. Through the alternating backpropagation, we obtain the gradients of each node’s output $E_{i}$ and its corresponding constants $w_{i}$, $b_{i}$.
Take a case as an example, as shown in Fig. 3c, the node $E_{1}=w_{1}f_{1}(E_{2},E_{5})+b_{1}$, where $f_{1}$ is the addition function, the gradients of $E_{1}$, $w_{1}$, and $b_{1}$ are:
$\displaystyle\nabla E_{1}$ $=\frac{\partial\mathcal{L}}{\partial E_{1}}\vspace{3ex}$ (5)
$\nabla w_{1}$ $=\frac{\partial\mathcal{L}}{\partial E_{1}}\frac{\partial E_{1}}{\partial w_{1}}=\nabla E_{1}\frac{\partial E_{1}}{\partial w_{1}}\vspace{3ex}$ (6)
$\nabla b_{1}$ $=\frac{\partial\mathcal{L}}{\partial E_{1}}\frac{\partial E_{1}}{\partial b_{1}}=\nabla E_{1}\frac{\partial E_{1}}{\partial b_{1}}=\nabla E_{1}\vspace{3ex}$ (7)
Where $\mathcal{L}$ is the loss function. The $E_{1}$ and constants are updated by the following formulas:
$\displaystyle E_{1_{\rm new}}=E_{1}-\alpha\nabla E_{1}$ (8)
$w_{1_{\rm new}}=w_{1}-\alpha\nabla w_{1}$ (9)
$b_{1_{\rm new}}=b_{1}-\alpha\nabla b_{1}$ (10)
Where $\alpha$ is the learning rate. According to the node chain rule, the gradients of $E_{2}$, $E_{5}$ are:
$\displaystyle\nabla E_{2}$ $=\frac{\partial\mathcal{L}}{\partial E_{2}}=\frac{\partial\mathcal{L}}{\partial E_{1}}\frac{\partial E_{1}}{\partial E_{2}}=\nabla E_{1}\frac{\partial E_{1}}{\partial E_{2}}\vspace{3ex}$ (11)
$\nabla E_{5}$ $=\frac{\partial\mathcal{L}}{\partial E_{5}}=\frac{\partial\mathcal{L}}{\partial E_{1}}\frac{\partial E_{1}}{\partial E_{2}}\frac{\partial E_{2}}{\partial E_{5}}=\nabla E_{2}\frac{\partial E_{2}}{\partial E_{5}}\vspace{3ex}$ (12)
For node 2, 5, $E_{2}=w_{2}(E3+E_{4})+b_{2}$, $E_{5}=w_{5}\sin(E_{6})+b_{5}$. $w_{2}$, $b_{2}$ are related to $E_{2}$, $W_{5}$, $b_{5}$ are related to $E_{5}$, thus, the gradients are calculated as:
$\displaystyle\nabla w_{2}$ $=\frac{\partial\mathcal{L}}{\partial w_{2}}=\frac{\partial\mathcal{L}}{\partial E_{1}}\frac{\partial E_{1}}{\partial E_{2}}\frac{\partial E_{2}}{\partial w_{2}}=\nabla E_{2}\frac{\partial E_{2}}{\partial w_{2}}\vspace{3ex}$ (13)
$\nabla b_{2}$ $=\frac{\partial\mathcal{L}}{\partial b_{2}}=\frac{\partial\mathcal{L}}{\partial E_{1}}\frac{\partial E_{1}}{\partial E_{2}}\frac{\partial E_{2}}{\partial b_{2}}=\nabla E_{2}\frac{\partial E_{2}}{\partial b_{2}}=\nabla E_{2}\vspace{3ex}$ (14)
$\nabla w_{5}$ $=\frac{\partial\mathcal{L}}{\partial w_{5}}=\frac{\partial\mathcal{L}}{\partial E_{1}}\frac{\partial E_{1}}{\partial E_{2}}\frac{\partial E_{2}}{\partial E_{5}}\frac{\partial E_{5}}{\partial w_{5}}=\nabla E_{5}\frac{\partial E_{5}}{\partial w_{5}}\vspace{3ex}$ (15)
$\nabla b_{5}$ $=\frac{\partial\mathcal{L}}{\partial b_{5}}=\frac{\partial\mathcal{L}}{\partial E_{1}}\frac{\partial E_{1}}{\partial E_{2}}\frac{\partial E_{2}}{\partial E_{5}}\frac{\partial E_{5}}{\partial b_{5}}=\nabla E_{5}\frac{\partial E_{5}}{\partial b_{5}}=\nabla E_{5}\vspace{3ex}$ (16)
Then, we update the values of the constants by subtracting the product of the learning rate and gradient.
1.3 Evaluate Activation Function
The updated output $E_{\rm new}$ of each neuron is arranged according to the pre-order traversal: $E_{\rm new}=[E_{1_{\rm new}},E_{2_{\rm new}},...,E_{m_{\rm new}}]$. We aim to select the activation from the library that makes the prediction more accurate. For the $i^{\rm th}$ node, $E_{i}=f_{i}(E^{i}_{\rm left_{old}},E^{i}_{\rm right_{old}})$, where $f_{i}$ is the addition function, $E^{i}_{\rm left_{old}}$ and $E^{i}_{\rm right_{old}}$ are the outputs of the left and right node of the $i^{\rm th}$ node, respectively. Suppose the values of the outputs are updated to $E_{i_{\rm new}}$, $E^{i}_{\rm left_{new}}$, and $E^{i}_{\rm left_{new}}$. The $E^{i}_{\rm left_{new}}$ and $E^{i}_{\rm left_{new}}$ are fed to the activation function library to calculate the new output of node $i$: $E_{i,j}^{\prime}=f_{i}(E^{i}_{\rm left_{new}},E^{i}_{\rm left_{new}})$, where $f_{i}\in\{+,-,\times,\div,sin,cos,log,sqrt,exp,relu,sigmoid,x\}$, $E_{i}^{\prime}\in\mathbb{R}^{12}$, $j=1,2,...,12$ is the index of the activation function library. Then we calculate the difference between $E_{i.j}^{\prime}$ and $E_{i_{\rm new}}$ and take the absolute value resulting in $G=\{g_{1},g_{2},...,g_{12}\}$, where $g_{j}=|E^{i}-E_{i_{\rm new}}|$. We select the activation function that has the smallest $g_{\rm min}$. To ensure steady training, we set a threshold (e.g. 0.01), $g_{\rm min}$ must lower than the threshold.
1.4 Substitute Activation Function
X-Net has diversified activation functions, and the exchange of activation functions brings difficulties to the forward propagation. For example, the arity of an activation function may change from one to two, the leaf node can not have a child node, etc. Therefore, we design several rules to ensure the X-Net can successfully conduct forward propagation after the activation functions have been updated, Fig. 2 shows the different types of substitution. Specifically, We set five rules:
  • (1)
    If the arity of the activation function does not change, substitute the old activation function to the new;
  • (2)
    If the activation function of a node changes from unary to binary, keep the left child and add a leaf node $x$ as the right child;
  • (3)
    If the activation function of a node changes from binary to unary, keep the child that has better fitting results and drop the other;
  • (4)
    If the activation function of a node changes to the leaf node, it does not have child nodes;
  • (5)
    If a leaf node changes to a unary node, the left child is added as the input of the node; If a leaf node changes to a binary node, the left and right children are added to the node.
1.5 Solutions for numerical computation
To avoid the gradient explosion problem, we revise some computation rules: (1) The outputs of all nodes are truncated during the computation process to prevent numerical overflow, specifically, $-V\leq E_{i}\leq V$, if $|E_{i}|\geq V$, $E_{i}=\frac{E_{i}}{|E_{i}|}V$; (2) We apply Gradient Clipping and limit the gradient in $[23,24]$. Assume the gradient value is $g_{i}$, which lies in the range $[-G_{max},G_{min}]\cup[G_{min},G_{max}]$ if $|g_{i}|\geq G_{max}$, then $g_{i}$ is set as $\frac{g_{i}}{|g_{i}|}G_{max}$; if $|g_{i}|\leq G_{min}$, then $g_{i}\frac{g_{i}}{|g_{i}|}G_{min}$; (3) We select the activation function in the basis of domain of definition, take the $\log(x)$ as an example, the $x$ must satisfy $x>0$.
@@ 1.6 Avoid Getting Stuck in Local Optimum
The network opts to fall into a local optimum in the process of optimization and training if no additional conditions are applied  (?). To prevent this problem from scratch, we set a count value $count$ as an indicator. If the loss remains unchanged or increases, $count=count+1$; If $count$ reaches a set threshold, we randomly select a node in X-Net and substitute a randomly selected activation function from the library. Through this “mutation” operation, X-Net is more likely to jump out of the local optimum. In our experiments, we found it was effective when the threshold was 20.
1.7 Adjust Learning Rate by Adaption Function: Ada-$\alpha$
The learning rate has a huge impact on the converging speed of the network and the search for the optimal solution  (?, ?, ?) in the training process. Therefore, we design an adaption function Ada-$\alpha$ which can adaptively and dynamically adjust the learning rate to achieve better fitting performance. Specifically, Ada-$\alpha$ is calculated by:
$\alpha=\frac{\tanh(e^{-|\mathcal{L}_{\rm pre}-\mathcal{L}_{\rm cur}|})}{a}.$ (17)
Where $\mathcal{L}_{\rm pre}$ and $\mathcal{L}_{\rm cur}$ represent the loss value in the previous iteration and current iteration, respectively. $a$ is the hyperparameter to adjust the range of learning rate.
Acknowledgments
Funding:
This work was supported in part by the National Natural Science Foundation of China under Grant 92370117, in part by CAS Project for Young Scientists in Basic Research under Grant YSBR-090, and in part by the Key Research Program of the Chinese Academy of Sciences under Grant XDPB22.
Authors contributions:
In this work, W.J. Li acted as the corresponding author, coordinating the research activities, injecting creativity into the project, and significantly enhancing its impact and scope. Y. Li was responsible for writing the paper, algorithm conception and design, and related experiments. L. Yu gave detailed guidance on the whole process of paper writing and experimental setup. M. Wu and L. Sun gave many useful suggestions for the design of the algorithm. J. Liu set up the structure of the article in detail. And carefully polished and improved the method part of the article. W.Q. Li and Y. Li collaborated on the writing of the abstract and introduction sections of the article. M. Hao, S. Wei, and Y. Deng helped the authors with some experiments. The completion of this work is inseparable from the efforts of each of the above. Thank you all for your help.
Competing interests:
All authors of the article have no competing interests
Supplementary materials
2 Pseudocode for X-Net
This is the pseudocode for the overall flow of our algorithm. We first initialize a tree neural network, and initialize the activation function $s_{i}$ for each node, and the parameters $[\mathcal{W},\mathcal{B}]$ of the network. Then we forward to compute the output of each neuron E. Next we compute the loss, and determine if the current loss on the validation set is due to the previous loss, if yes, we save the current network architecture and the optimal $R^{2}$. Train the network and repeat until a predetermined number of iterations is reached.